Instructions : Answer true or false and explain the answer

(1) Suppose s = 0.15, Y = 4200, K = 6100, n = 0.03, g=0.03 and δ= 0.10. This makes national saving smaller than steady-state investment, so that the amount of capital per effective worker will be falling.

(2) In the Solow growth model, given the values of A=120, s=0.11, n=0.03, g=0.07, and δ=0.08, the economy has an equilibrium growth rate of real GDP per effective labour, (Y/AL), equal to 8 percent.

(3) In the graph of the Solow growth model, at any point to the left of the steady-state intersection we have national saving per effective labour greater than steady-state investment per person, causing (K/AL) to increase.

(4) In the Solow growth model, an increase in the marginal propensity to consume shifts the steady-state investment line downward with the implied change in the capital stock resulting in a higher standard of living in the long run.
1
Expert's answer
2021-08-04T09:37:35-0400

Solution:

1.). The statement is false.

The formula for k is as follows:

K = ("\frac{s}{n +δ} )^{2}"


K = ("\frac{0.15}{0.03 +0.10} )^{2}"


K = ("\frac{0.15}{0.13} )^{2}"


K = "\frac{0.0225}{0.169} = 1.33"


As per the above calculation, K is positive which means that the amount of capital per effective worker will be increasing.


2.). The statement is false:

Y/AL = (n+g+δ)k

Y/AL = (0.03+0.07+0.08)"\times"120

Y/AL = 0.18 "\times"120 = 22"\%"


3.). The statement is false:

Solow model does not reflect national saving per effective labor. It rather shows the changes in the output level in an economy over a particular period of time.


4.). The statement is correct:

MPC refers to that part of consumption that a consumer purchases or utilizes due to an increase in income, instead of saving. When a person’s income is higher, the MPC will decrease, whereas when a person’s income is low, the MPC will be much higher. MPC tends to be high when we have low-income earners since most of the person’s income has to be allocated to subsistence consumption.